| x| >1;let y =1+x+x^2+…………to infinity then dy/dx equal to


 

Grade:

2 Answers

student iit
14 years ago

This is infinite GP where common ration x>1.

S(n)= 1(x^n-1)/(x-1)= x^n-1/x-1

Now use d(u/v)/dx formula to find the drivative.

Badiuddin askIITians.ismu Expert
14 years ago

hi nitish kumar

| x| >1;

 

 y =1+x+x^2+…………to infinity

 

dy/dx =1+2x+3x2+4x3  .....................infinity

 

clearl for x>1  dy/dx tends to +infinity

 

  and for x<-1  dy/dx  tends to either + infinity or - infinity

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation. All the best nitish
 
 

Regards,

Askiitians Experts

Badiuddin